OpenStudy (anonymous):

can someone please explain to me how you would get 5/8 from the problem 1/2 + 1/8 just getting back to school havent been in school seen 2000

OpenStudy (anonymous):

1/2+1/8=? for adding these first we need to make the denominators same....to do so take the l.c.m..here the lcm is 8.multiply 1/2 by 4,then it becoms 4/8.now the denominators are same,so can proceed with the addition..ie,4/8+1/8=5/8..

OpenStudy (anonymous):

in this problem,1/2 can also be written as 4/8 since 4 is half of 8 as 1 is half of 2. 4/8+1/8 since the denominators(the numbers at the bottom) are equal the numerators(the numbers on top )can be added therefore 4/8+1/8=5/8 hope this helps